OpenStudy (anonymous):

Show all work in dividing the quantity y-squared minus 3y plus 2 all over the quantity y-squared plus 5y plus 6 by the quantity y-squared plus 2y minus 8 all over the quantity y-squared plus y minus 12

OpenStudy (anonymous):

\[\left(\begin{matrix}y^2-3y+2 \\ y^2+5y+6\end{matrix}\right)\div \left(\begin{matrix}y^2+2y-8 \\ y^2+y-12\end{matrix}\right)\]

OpenStudy (anonymous):

Can someone help me? It's urgent, I'm on a timed test. Please and thank you.